OpenStudy (anonymous):

how would you fine the derivative of this please: f(x) = (−9 x^2 + 8)^5 (4 x^2 + 2)^9

OpenStudy (anonymous):

looks like a product rule combined with a chain rule problem

myininaya (myininaya):

use product rule+chain rule

myininaya (myininaya):

\[[(-9x^2+8)^5]'(4x^2+2)^9+(-9x^2+8)^5[(4x^2+2)^9]'\]

OpenStudy (anonymous):

use \[(gh)'=g'h+h'g\] with \[g(x)=(-9x^2+7)^5,g'(x)=5(-9x^2+8)^6\times -18x\] \[h(x)=(4x^2+2)^9,h'(x)=9(4x^2+2)^8\times 8x\]

OpenStudy (anonymous):

i lose

OpenStudy (anonymous):

time to go to work in any case

OpenStudy (anonymous):

was there a typo?

OpenStudy (anonymous):

what is a minus sign between friends? later

myininaya (myininaya):

peace i need to read a paper